# AI for Just Work: Constructing Diverse Imaginations of AI beyond “Replacing Humans”

### v1

@@ Abstract
The AI community usually focuses on “how” to develop AI techniques, but lacks thorough open discussions on “why” we develop AI. Lacking critical reflections on the general visions and purposes of AI may make the community vulnerable to manipulation. In this position paper, we explore the “why” question of AI. We denote answers to the “why” question the imaginations of AI, which depict our general visions, frames, and mindsets for the prospects of AI. We identify that the prevailing vision in the AI community is largely a monoculture that emphasizes objectives such as replacing humans and improving productivity. Our critical examination of this mainstream imagination highlights its underpinning and potentially unjust assumptions. We then call to diversify our collective imaginations of AI, embedding ethical assumptions from the outset in the imaginations of AI. To facilitate the community’s pursuit of diverse imaginations, we demonstrate one process for constructing a new imagination of “AI for just work,” and showcase its application in the medical image synthesis task to make it more ethical. We hope this work will help the AI community to open dialogues with civil society on the visions and purposes of AI, and inspire more technical works and advocacy in pursuit of diverse and ethical imaginations to restore the value of AI for the public good.
AI Ethics, AI Research Priorities, Imagination, Machine Learning
@@ 1 The imagination of AI: What is AI for?
The state-of-the-art data-driven technologies, now primarily called artificial intelligence/machine learning (AI/ML) (Goodlad, 2023), have gained great momentum and societal attention in the past decade. In the AI community, we often ask “how” to propose, develop, and deploy AI techniques, but rarely reflect on the “why” question on the purposes of AI. Lacking critical reflections on the fundamental “why” question may render the community susceptible to being trapped by suboptimal objectives, influenced by distorted research agenda, and manipulated for unjust purposes (Greenhalgh, 2024; Brynjolfsson, 2022; Sismondo, 2018). In this work, we explore the “why” question in AI. We ask, “Why do we develop AI? What purposes does AI serve?” We denote answers to these inquiries the imaginations of AI, which depict the overarching motivations of AI, i.e., how AI can be helpful to humanity and the planet. Extended from the concept of imaginary in social science (Benjamin, 2024; Earle, 2020), the imagination of AI is the general vision, frame, and mindset we have for the prospects of AI technologies. The imagination of AI sets the agenda on what are important goals to pursue in AI that reflect our value systems, and shapes assumptions, design choices, and objectives when developing specific AI techniques.
@@
Of course, different people have different imaginations of how AI can be helpful for humanity and the planet. As a community, our collective imaginations should also be diversified to reflect and accommodate a variety of pursuits from people with different backgrounds, values, and interests. While different visions of AI exist in the community, and such works usually appear in venues and subfields such as AI for (Social) Good, AI for Science, or the ICML Position Paper Track, they are either at the inception or in a relatively niche position regarding their overall influence [^1] There are more critiques with different visions of AI from other fields like critical theory, FAccT, STS, see Appendix B for details. Our work draws on these critiques and diagnoses, and contributes to the AI community by adapting them to the technical context to bring them into the community’s focus, and bridging diagnoses with treatment in AI technical practice.. In comparison, the long-established mainstream imaginations of AI in the community are dominated by a monoculture that is shaped by “a small and largely homogeneous group of people” (Burrell & Metcalf, 2024). This mainstream imagination of AI emphasizes progress and performance, encoding values such as performance, generalization, efficiency, and novelty according to a study on 100 highly-cited papers in ICML and NeurIPS (Birhane et al., 2022). The prominent manifestation of this imagination within the AI community and in the public is the AI objective to automate human occupations and to surpass and replace humans, evidenced by a recent large-scale survey with 2,778 AI researchers on the future of AI (Grace et al., 2024) and the widespread societal anxiety of being replaced by AI (Ivanov et al., 2020).
While the mainstream imagination has some merits, it is not without flaws. In this imagination, there is no room to encode values that can directly improve human and environmental welfare at the outset in the blueprint of AI, such as justice, democracy, and sustainability. Instead, these ethical values are regarded as secondary priority and mainly used as post-harm fix rather than prevention from the root. Furthermore, fixating on a single-minded imagination is a problem for the community, indicating other possible imaginations and the diverse values underpin them are suppressed (Marcus, 2024; Burrell & Metcalf, 2024; Rudin & Wagstaff, 2013). People who envision different imaginations could feel unwelcome and their contributions unrecognized, which is the partial reason for the current lack of diversity and inclusion in the AI community (Wajcman & Young, 2023).
@@
To address these issues, in this work we call for 1) diversified collective imaginations of AI that 2) synergize both ethical values and AI development (Fig. 1). We first synthesize multidisciplinary evidence and rationales to understand why the current mainstream imagination itself is largely unethical and can cause harms (Section 2). We then highlight the important role of imaginations in technical and social actions (Section 3), and demonstrate a process of constructing a new imagination of “AI for just work” that bakes ethical values in its grounding assumptions and technical practice (Section 4). We hope this work can stimulate the AI community to open public debates and discussions with civil society on the “why” question of AI, especially with those who are vulnerable, marginalized, or negatively affected by AI, and support diverse imaginations, values, and works for the ethical pursuit of AI development, such that to restore the value of AI technology as a public good.
2 The mainstream imagination of AI (MIA) is not as ethical as we assume
2.1 Alternative views: MIA and its justifying statements
The mainstream imagination of AI (MIA) is self-explanatory and seldom questioned, probably because it is based on fundamental assumptions that are the basic beliefs within the AI community and in society. The typical logic to justify that MIA is an ethical pursuit for social good can be summarized in the following statements (S1-S4) [^2] We note of course that even for the mainstream viewpoints in the AI community, there are substantial diversity and nuances in viewpoints that S1-S4 do not capture.:
@@ 2.1 Alternative views: MIA and its justifying statements
Given the critical role of these statements in legitimizing the ethical pursuits of MIA and their consequences in society, these statements cannot be taken at face value. Rather, the AI community has the responsibility to either critically examine the validity and ethics of these statements, or provide solid evidence and rationales to defend these statements in the face of counterarguments to establish their validity and ethics in supporting MIA. Since both aspects are currently missing in the community, next we conduct a critical examination on MIA and its justifying statements S1-S4 by synthesizing evidence and rationales from philosophy of science (S1), ethics (S2), feminism (S2), economics (S3), and social science (S4).
2.2 Critical examination of MIA
S1. AI can be omniscient.
The imagination to create a God-like, super human intelligence that is capable of solving any hard problem for humanity at our will incarnates the unspoken monotheistic mindset underlying modern science and technology. Frank et al. (2024) name this mindset the blind spot worldview. It is characterized by the philosophical perspective that there exists a rational, “disembodied, God’s-eye perspective” for us to gain “access to a perfectly knowable, timeless objective reality” (Frank et al., 2024). This worldview divorces the knower and the known, subjectivity and objectivity, forgets their mutual relationship, and unnecessarily considers objectivity to have a higher hierarchy than subjectivity regarding their closeness to the idealized, God’s-eye view of truth (Fig. 2). Entangled with culture, politics, and economy, this limited perspective on science and technology was woven into the fabric of our value systems in society, which values rationality, abstraction, objectification, quantification, order, control, and downgrades experience, intuition, sense, emotion, fluidity, and ambiguity. It, in accordance, frames how we regard the human experience and the natural world, which relegates them to pure resources to control and exploit (Frank et al., 2024). As Frank et al. (2024) argue, this “impoverishes the living world and our experience” and leads to existential crises such as environmental destruction, global pandemic, and algorithmic surveillance. Next, we examine the particular costs and harms of MIA to human labor and society.
Figure 2: Visualization of two contrasting worldviews of science and technology including computational models. (A) The normal worldview: it acknowledges and values the fact that computational models are built upon human experience (embodied as labor, knowledge, dataset, infrastructure, etc.), and rely on it to validate and apply models in the real world; human experience and models are like pillars and floor in the figure: they are mutually dependent on each other to better understand our living world; computational models are useful reductive abstractions of, but cannot fully represent, the complex real-world phenomena they model (the well-known aphorism that “all models are wrong, but some are useful” (Box, 1976)). (B) The blind spot worldview: it wrongly assumes that computational models are dominant over human experience, “belies AI’s fundamental dependence on physical nature and socially organized, collective human knowledge” (“the amnesia of experience”), “substitut[es] computation for genuine embodied intelligence,” and “treat[s] abstract computational models as if they were concretely real” (“confuse the map with the territory”). Unspecified quotations in the figure and caption are from Frank et al. (2024).
S2. Replacing human with AI can free human.
The imagination to replace humans by AI forgets the mutually dependent relationship between computational models and human experience (Fig. 2-A), and assumes that AI can be dominant over human experience, which is a manifestation of the blind spot worldview (Fig. 2-B). By forgetting humans’ fundamental role as the “epistemic footing” (Frank et al., 2024) for AI, it 1) “wrongs someone in their capacity as a subject of knowledge, and thus in a capacity essential to human value,” which manifests epistemic injustice in ethics (Fricker, 2007). It also 2) impoverishes our human experience of work that is supposed to be “meaningful, fulfilling activities” as the “social means necessary to live a flourishing life” (Wright, 2021). We expand the two interrelated harms below.
First, the imagination of labor replacement for human liberation is not new. Atanasoski & Vora (2019b) identify that the narrative of AI replacement shares similar patterns and assumptions with the long-standing labor replacement phenomena since the inception of capitalism: “the emergence of reproductive unpaid women’s work in Europe [due to witch hunts]…replace[d] the rebellious peasants opposing the enclosure of common lands and resources;” In the US plantations, farmers and indentured servants were replaced by slaves; In globalization, domestic or in-house workers were replaced by offshore or outsourced workers; Nowadays in gig economy, full-time, permanent workers are replaced by precarious workers such as contract, part-time, or gig workers, who in turn generate data to train AI and are expected to be replaced by AI, a.k.a., the machine slaves [^3] The word “robot” is from Slavic linguistic root “rab,” which means “slave” (Jordan, 2019). (Atanasoski & Vora, 2019b). In each labor replacement, the original workers are replaced by workers who are less paid, poorly treated, enjoy much less freedom in their work, are less likely to rebel, are atomized and isolated thus harder to unionize and have little to no bargaining power. In other words, labor replacements greatly reduce the power of workers who are replaced and who replace others; and the proposed human liberation from labor replacements is merely “enabling and improving the lives of privileged subjects” (Atanasoski & Vora, 2019b).
For such labor exploitation and oppression to happen, systematic violence (Federici, 2004), state and imperial forces (Mies, 1986), and divide and conquer strategy are needed to establish gendered, racialized, and other social hierarchies that legitimize the exploitation of the “degraded and devalued others—those who were never fully human” (Atanasoski & Vora, 2019b). The MIA to replace humans is no different: by assuming a new human-machine hierarchy, it still reproduces the same logic of racialized and gendered social hierarchies that continue to legitimize more hidden labor exploitation (Atanasoski & Vora, 2019b). Specifically, as we mentioned, there are two groups of workers who are prominently harmed by AI replacement: those whose jobs are taken by AI, and those who work behind or with AI, such as Uber drivers (Rosenblat, 2018), Amazon Mechanical Turk workers (Gray & Suri, 2019), and Amazon warehouse workers (MacGillis, 2021). The latter group suffers from more hidden harms and exploitation, because their work and contributions are often purposely made invisible and are attributed as the “intelligence” of the machine (Atanasoski & Vora, 2019b), reflecting the blind spot worldview and epistemic injustice. Since AI constantly depends on human labor and experience (Frank et al., 2024), as long as the hierarchical assumption in AI exists, these harms will persist no matter how AI advances.
Second, the MIA to replace humans not only intensified labor exploitation as described above, but also impoverished the concept of work. Ironically, in the discourse of AI replacing humans, the bogus “autonomous” aspect of machine is emphasized, but the real human autonomy, which is one of the five key principles in AI ethics (Floridi et al., 2018), is ignored. Consequently, workers and work tend to be organized around the capabilities and limitations of the machine, not the other way around (Delfanti, 2021). As Frank et al. (2024) note, one aspect of “the computational blind spot…is failing to see how we are led to remake the world so that it gears into the limitations of our computational systems.” The deprived workers’ power and autonomy reduced work to a single dimension of the necessary means (or drudgery) to earn one’s living (Resnikoff, 2021), whereas the multifaceted meaning of work, such as a social means to live a meaningful and dignified life or connecting with people and the community, can only be fully enjoyed by a privileged few. The scope of work was also narrowed that defines whose and what kinds of activities count as work and can enjoy corresponding protection, support, and recognition, and what don’t (Mies, 1986). The refashion of our living world and work to adapt to machine greatly limits everyone’s living experience and options on how we can live a flourishing life in a originally meaning-rich world.
S3. AI can promote productivity and prosperity.
Although S2 has many unreasonable aspects and obvious harms including the above counterarguments, a strong support for it is in S3 that AI advances can raise productivity, which will eventually benefit us all despite the harms. However, economic studies have shown that such “lift all boats” effect does not happen automatically with technological progress, but is the result of both 1) our choices of how technology is developed and 2) active labor struggles and bargaining (Acemoglu & Johnson, 2023).
First, AI does not necessarily lead to productivity gain, and not any types of productivity gain can precondition the broad-based prosperity. Not all AI technologies increase productivity. For example, AI-enabled surveillance techniques are not mainly focused on improving productivity, but to impose more control over workers, reduce their power, and save on wage costs (Acemoglu & Johnson, 2023). Furthermore, replacing labor with AI tends to shift the production in favor of capital and against labor, as it creates a strong displacement effect that reduces labor demand and wages, which is in the opposite direction to broad-based prosperity. When the productivity gain from automation is not enough to offset the displacement effect, it is the so-so automation or excessive automation that lowers productivity growth and widens inequality  (Acemoglu & Restrepo, 2019a, c). Empirical evidence in the US and Germany shows that exposure to industrial robots or automation reduced labor income share, which impacted low-skill workers more and thus exacerbated inequality (Acemoglu & Restrepo, 2022, 2020; Dauth et al., 2017). The displacement effect may be countervailed and labor share may be increased only when workers’ marginal productivity is increased. Distinct from the standard definition of productivity (or average productivity), which is a worker’s average output, marginal productivity is “the additional contribution that one more worker brings by increasing production or by serving more customers” (Acemoglu & Johnson, 2023). Marginal productivity comes from the increased productivity that is large enough to expand jobs and employment, from technologies that can aid workers, and more importantly, from technologies that can create new labor-intensive tasks “in which labor has a comparative advantage” (Acemoglu & Restrepo, 2019b, c). These directions, as Acemoglu & Johnson (2023) point out, are all highly dependent on the different visions we choose to develop AI.
Second, the increased marginal productivity alone is not enough for broad-based prosperity, but should be accompanied by strong labor power that guides the direction of technological development and ensures that marginal productivity is shared with workers (Acemoglu & Johnson, 2023). The increased demand for labor does not automatically increase workers’ wages, because employer and employee are in a coercive relationship, the employer may not face external competitors for workers, and wages are determined not only by market forces but also by bargaining power (Acemoglu & Johnson, 2023). Acemoglu & Johnson (2023) review the economic history and technical change from the Middle Age to present, and associate the changes in workers’ wages, welfare, working conditions, and directions of technical development with changes in social and political power.
S4. Harms of AI are due to malicious use.
@@ 2.2 Critical examination of MIA
Our examination of S1-S3 shows that unethical issues in AI are rooted in intertwined causes encoded in the MIA including: a limited philosophical worldview, unjust hierarchy in social structures of labor, and imbalanced economic and political power. Embedded in the social system, AI technology acts as a mediator, is shaped by these root causes and in turn shapes the unethical impacts of AI in society (Matthewman, 2011; Boenig-Liptsin et al., 2022). Tackling the malicious use to solve unethical issues in AI only treats the symptoms, not the disease. Similarly, the imagination that AI can fix its own issues in S1 is not based on valid logic. Solving a problem requires first conducting a precise diagnosis of the root cause, and then identifying treatment that can target the diagnosis. There is no guarantee that the right treatment will always fall into the realm of the technical toolbox, and claiming AI can fix any issues is like selling a cure-all elixir (Narayanan & Kapoor, 2024).

### v2

@@ Abstract
The AI community usually focuses on “how” to develop AI techniques, but lacks thorough open discussions on “why” we develop AI. Lacking critical reflections on the general visions and purposes of AI may make the community vulnerable to manipulation. In this position paper, we explore the “why” question of AI. We denote answers to the “why” question the imaginations of AI, which depict our general visions, frames, and mindsets for the prospects of AI. We identify that the prevailing vision in the AI community is largely a monoculture that emphasizes objectives such as replacing humans and improving productivity. Our critical examination of this mainstream imagination highlights its underpinning and potentially unjust assumptions. We then call to diversify our collective imaginations of AI, embedding ethical assumptions from the outset in the imaginations of AI. To facilitate the community’s pursuit of diverse imaginations, we demonstrate one process for constructing a new imagination of “AI for just work,” and showcase its application in the medical image synthesis task to make it more ethical. We hope this work will help the AI community to open critical dialogues with civil society on the visions and purposes of AI, and inspire more technical works and advocacy in pursuit of diverse and ethical imaginations to restore the value of AI for the public good.
@@ 1 The imagination of AI: What is AI for?
The state-of-the-art data-driven technologies, now primarily called artificial intelligence/machine learning (AI/ML) [53], have gained great momentum and societal attention in the past decade. In the AI community, we often ask “how” to propose, develop, and deploy AI techniques, but rarely reflect on the “why” question on the purposes of AI. Lacking critical reflections on the fundamental “why” question may render the community susceptible to being trapped by suboptimal objectives, influenced by distorted research agenda, and manipulated for unjust purposes [23, 57, 28, 100]. In this work, we explore the “why” question in AI. We ask, “Why do we develop AI? What purposes does AI serve?” We denote answers to these inquiries the imaginations of AI, which depict the overarching motivations of AI, i.e., how AI can be helpful to humanity and the planet. Extended from the concept of imaginary in social science [18, 42], the imagination of AI is the general vision, frame, and mindset we have for the prospects of AI technologies. The imagination of AI sets the agenda on what are important goals to pursue in AI that reflect our value systems, and shapes assumptions, design choices, and objectives when developing specific AI techniques.
Of course, different people have different imaginations of how AI can be helpful for humanity and the planet. As a community, our collective imaginations should also be diversified to reflect and accommodate a variety of pursuits from people with different backgrounds, values, and interests. While AI community has different visions of AI, and such visions and works usually appear in venues and subfields such as AI for (Social) Good, AI for Science, or FAccT (Fairness, Accountability, and Transparency), they are either in the inception or in a relatively niche position regarding their overall influence [^1] There are more critiques with different visions of AI from other fields including critical theory and STS (Science, Technology and Society), see Appendix D for details. Our work draws on these critiques and diagnoses, and contributes to the AI community by adapting them to the technical context to bring them into the community’s focus, and bridging diagnoses with treatment in AI technical practice.. In comparison, the long-established mainstream imaginations of AI in the community are dominated by a monoculture [23, 67, 75] that is shaped by “a small and largely homogeneous group of people” [31]. This mainstream imagination of AI emphasizes progress and performance, encoding values such as performance, generalization, efficiency, and novelty according to a study on 100 highly-cited papers in ICML and NeurIPS [20]. The prominent manifestation of this imagination within the AI community and in the public is the AI objective to automate human occupations and to surpass and replace humans, evidenced by a recent large-scale survey with 2,778 AI researchers on the future of AI [54] and the widespread societal anxiety of being replaced by AI [60].
While the mainstream imagination has some merits, it is not without flaws. In this imagination, there is no room to encode values that can directly improve human and environmental welfare at the outset in the blueprint of AI, such as justice, democracy, and sustainability. Instead, these ethical values are regarded as secondary priority and mainly used as post-harm fix rather than prevention from the root. Furthermore, fixating on a single-minded imagination is a problem for the community, indicating other possible imaginations and the diverse values underpin them are suppressed [73, 31, 97]. People who envision different imaginations could feel unwelcome and their contributions unrecognized, which is the partial reason for the current lack of diversity and inclusion in the AI community [105].
@@
To address these issues, in this work we call for 1) diversified collective imaginations of AI that 2) synergize both ethical values and AI development (Fig. 1). We first synthesize multidisciplinary evidence and rationales to understand why the current mainstream imagination itself is largely unethical and can cause harms (Section 2). We then highlight the important role of imaginations in technical and social actions (Section 3), and demonstrate a process of constructing a new imagination of “AI for just work” that bakes ethical values in its grounding assumptions and technical practice (Section 4). We hope this work can stimulate the AI community to open public debates and discussions with civil society on the “why” question of AI, especially with those who are vulnerable, marginalized, or negatively affected by AI, and support diverse imaginations, values, and works for the more ethical pursuit of AI development, such that to restore the value of AI technology as a public good.
2 The mainstream imagination of AI is not as ethical as we assume
2.1 Alternative views: the mainstream imagination of AI (MIA)
As mentioned in Section 1, although individuals in the AI community have different imaginations of AI, such tendency may not proportionately propagate to the community level, as shown by empirical evidence [20, 54, 60]. This non-linear relationship that “community $\neq$ a collection of individuals” indicates there are latent factors that influence how individuals’ imaginations and works are selected, recognized, and supported, usually through processes such as peer review, funding, and career advancement. These factors, termed as social structures [^2] We provide more background information on the concept of social structure in Section D.1.2, which clarifies that a community is not just individuals but also consists of social structures. Understanding the role of social structures is vital to avoid taking critiques on structural issues personally, and enables us to identify and address root causes in structural issues, which in turn can better support individuals’ pursuits for ethical AI., pave the paths of least resistance [63] and in turn influence individuals’ framing of their work. Our discussion focuses on the community’ collective imaginations, which manifests features of the social structures in the AI community and its embedded society, not features of the individuals. The community’s collective imaginations are dominated by a monoculture, which we name the mainstream imagination of AI (MIA). MIA emphasizes notions of an idealized picture of technological progress with unscoped technical capabilities (S1), human-level performance, automation, human replacement (S2), productivity improvement (S3), and attributing AI risks and harms largely to mistakes and malicious use (S4). These notions are then associated with the anticipated benefits to humanity and the planet to justify the legitimacy and ethics of MIA. A typical logic of justification is abstracted in statements [^3] Although we try to faithfully abstract the typical ideas and justifications of MIA into S1-S4 based on the cited references in Section 2.2, these statements cannot fully capture the substantial variations of the community’s ideas. This is an intrinsic limitation of our analysis. S1-S4 in Section 2.2.
Given the critical role of these statements in legitimizing the ethical pursuits of MIA and their consequences in society, these statements cannot be taken at face value. Rather, the AI community has the responsibility to either critically examine the validity and ethics of these statements, or provide solid evidence and rationales to defend these statements in the face of counterarguments to establish their validity and ethics in supporting MIA. Since both aspects are currently missing in the community, a critical examination on MIA and its justifying statements S1-S4 would be timely and beneficial to the healthy development of AI and the community. Next, we conduct a critical examination of MIA. Since MIA uses justifications related to topics of the worldviews of technology, ethics, labor, society, and economic objectives, our examination will also synthesize knowledge from these disciplines [^4] Specifically, we synthesize evidence and rationales from philosophy of science (S1), ethics (S2), feminism (S2), economics (S3), and social science (S4), and offers simplified introductory discussions on these rich and complex disciplines. Each discipline has various and usually competing theories and standpoints on the given topics, and our arguments prioritize perspectives that represent the vulnerable, marginalized, and disproportionately impacted groups over the majority [19]. Thus, our arguments are limited by the particular ethical and philosophical standpoint we choose. The limitation in our discussion opens rooms for readers to discuss from other perspectives.. To facilitate readers’ understanding and discussions of these interdisciplinary topics, we supplement our discussion with a glossary table in Appendix C, and a bibliographic essay on the multidisciplinary knowledge background in Appendix D.
2.2 Critical examination of MIA
@@
Figure 2: Visualization of two contrasting worldviews of science and technology including computational models. (A) The normal worldview: it acknowledges and values the fact that computational models are built upon human experience (embodied as labor, knowledge, dataset, infrastructure, etc.), and rely on it to validate and apply models in the real world; human experience and models are like pillars and floor in the figure: they are mutually dependent on each other to better understand our living world; computational models are useful reductive abstractions of, but cannot fully represent, the complex real-world phenomena they model (the well-known aphorism that “all models are wrong, but some are useful” [25]). (B) The blind spot worldview: it wrongly assumes that computational models are dominant over human experience, “belies AI’s fundamental dependence on physical nature and socially organized, collective human knowledge” (“the amnesia of experience”), “substitut[es] computation for genuine embodied intelligence,” and “treat[s] abstract computational models as if they were concretely real” (“confuse the map with the territory”). Unspecified quotations in the figure and caption are from Frank et al., [47].
The imagination to create a God-like, super human intelligence that is capable of solving hard problems [82, 65] for humanity at our will incarnates the unspoken monotheistic mindset underlying modern science and technology. Frank et al., [47] name this mindset the blind spot worldview. It is characterized by the philosophical perspective that there exists a rational, “disembodied, God’s-eye perspective” for us to gain “access to a perfectly knowable, timeless objective reality” [47]. This worldview divorces the knower and the known, subjectivity and objectivity, forgets their mutual relationship, and unnecessarily considers objectivity to have a higher hierarchy than subjectivity regarding their closeness to the idealized, God’s-eye view of truth (Fig. 2). Entangled with culture, politics, and economy, this limited perspective on science and technology was woven into the fabric of our value systems in society, which values rationality, abstraction, objectification, quantification, order, control, and downgrades experience, intuition, sense, emotion, fluidity, and ambiguity. It, in accordance, frames how we regard the human experience and the natural world, which relegates them to pure resources to control and exploit [47]. As Frank et al., [47] argue, this “impoverishes the living world and our experience” and leads to existential crises such as environmental destruction, global pandemic, and algorithmic surveillance. Next, we examine the particular costs and harms of MIA to human labor and society.
@@
The imagination to replace humans by AI forgets the mutually dependent relationship between computational models and human experience (Fig. 2-A), and assumes that AI can be dominant over human experience, which is a manifestation of the blind spot worldview (Fig. 2-B). By forgetting humans’ fundamental role as the “epistemic footing” [47] for AI, it 1) “wrongs someone in their capacity as a subject of knowledge, and thus in a capacity essential to human value,” which manifests epistemic injustice in ethics [50]. It also 2) impoverishes our human experience of work that is supposed to be “meaningful, fulfilling activities” as the “social means necessary to live a flourishing life” [110]. We expand the two interrelated harms below.
First, the imagination of labor replacement for human liberation is not new. Atanasoski and Vora, 2019b [15] identify that the narrative of AI replacement shares similar patterns and assumptions with the long-standing labor replacement phenomena since the inception of capitalism: “the emergence of reproductive unpaid women’s work in Europe [due to witch hunts]…replace[d] the rebellious peasants opposing the enclosure of common lands and resources;” In the US plantations, farmers and indentured servants were replaced by slaves; In globalization, domestic or in-house workers were replaced by offshore or outsourced workers; Nowadays in gig economy, full-time, permanent workers are replaced by precarious workers such as contract, part-time, or gig workers, who in turn generate data to train AI and are expected to be replaced by AI, a.k.a., the machine slaves [^5] The word “robot” is from Slavic linguistic root “rab,” which means “slave” [64]. [15]. In each labor replacement, the original workers are replaced by workers who are less paid, poorly treated, enjoy much less freedom in their work, are less likely to rebel, are atomized and isolated thus harder to unionize and have little to no bargaining power. In other words, labor replacements greatly reduce the power of workers who are replaced and who replace others; and the proposed human liberation from labor replacements is merely “enabling and improving the lives of privileged subjects” [15].
For such labor exploitation and oppression to happen, systematic violence [44], state and imperial forces [77], and divide and conquer strategy are needed to establish gendered, racialized, and other social hierarchies that legitimize the exploitation of the “degraded and devalued others—those who were never fully human” [15]. The MIA to replace humans is no different: by assuming a new human-machine hierarchy, it still reproduces the same logic of racialized and gendered social hierarchies that continue to legitimize more hidden labor exploitation [15]. Specifically, as we mentioned, there are two groups of workers who are prominently harmed by AI replacement: those whose jobs are taken by AI, and those who work behind or with AI, such as Uber drivers [96], Amazon Mechanical Turk workers [56], and Amazon warehouse workers [72]. The latter group suffers from more hidden harms and exploitation, because their work and contributions are often purposely made invisible and are attributed as the “intelligence” of the machine [15], reflecting the blind spot worldview and epistemic injustice. Since AI constantly depends on human labor and experience [47], as long as the hierarchical assumption in AI exists, these harms will persist no matter how AI advances.
Second, the MIA to replace humans not only intensified labor exploitation as described above, but also impoverished the concept of work. Ironically, in the discourse of AI replacing humans, the bogus “autonomous” aspect of machine is emphasized, but the real human autonomy, which is one of the five key principles in AI ethics [45], is ignored. Consequently, workers and work tend to be organized around the capabilities and limitations of the machine, not the other way around [40]. As Frank et al., [47] note, one aspect of “the computational blind spot…is failing to see how we are led to remake the world so that it gears into the limitations of our computational systems.” The deprived workers’ power and autonomy reduced work to a single dimension of the necessary means (or drudgery) to earn one’s living [93], whereas the multifaceted meaning of work, such as a social means to live a meaningful and dignified life or connecting with people and the community, can only be fully enjoyed by a privileged few. The scope of work was also narrowed that defines whose and what kinds of activities count as work and can enjoy corresponding protection, support, and recognition, and what don’t [77]. The refashion of our living world and work to adapt to machine greatly limits everyone’s living experience and options on how we can live a flourishing life in a originally meaning-rich world.
@@
Although S2 has many unreasonable aspects and obvious harms including the above counterarguments, a strong support for it is in S3 that AI advances can raise productivity, which will eventually benefit us all despite the harms. However, economic studies have shown that such “lift all boats” effect does not happen automatically with technological progress, but is the result of both 1) our choices of how technology is developed and 2) active labor struggles and bargaining [2].
First, AI does not necessarily lead to productivity gain, and not any types of productivity gain can precondition the broad-based prosperity. Not all AI technologies increase productivity. For example, AI-enabled surveillance techniques are not mainly focused on improving productivity, but to impose more control over workers, reduce their power, and save on wage costs [2]. Furthermore, replacing labor with AI tends to shift the production in favor of capital and against labor, as it creates a strong displacement effect that reduces labor demand and wages, which is in the opposite direction to broad-based prosperity. When the productivity gain from automation is not enough to offset the displacement effect, it is the so-so automation or excessive automation that lowers productivity growth and widens inequality  [6, 4]. Empirical evidence in the US and Germany shows that exposure to industrial robots or automation reduced labor income share, which impacted low-skill workers more and thus exacerbated inequality [8, 7, 39]. The displacement effect may be countervailed and labor share may be increased only when workers’ marginal productivity is increased. Distinct from the standard definition of productivity (or average productivity), which is a worker’s average output, marginal productivity is “the additional contribution that one more worker brings by increasing production or by serving more customers” [2]. Marginal productivity comes from the increased productivity that is large enough to expand jobs and employment, from technologies that can aid workers, and more importantly, from technologies that can create new labor-intensive tasks “in which labor has a comparative advantage” [5, 4]. These directions, as Acemoglu and Johnson, [2] point out, are all highly dependent on the different visions we choose to develop AI.
Second, the increased marginal productivity alone is not enough for broad-based prosperity, but should be accompanied by strong labor power that guides the direction of technological development and ensures that marginal productivity is shared with workers [2]. The increased demand for labor does not automatically increase workers’ wages, because employer and employee are in a coercive relationship, the employer may not face external competitors for workers, and wages are determined not only by market forces but also by bargaining power [2]. Acemoglu and Johnson, [2] review the economic history and technical change from the Middle Age to present, and associate the changes in workers’ wages, welfare, working conditions, and directions of technical development with changes in social and political power.
@@
Our examination of S1-S3 shows that unethical issues in AI are rooted in intertwined causes encoded in the MIA including: a limited philosophical worldview, unjust hierarchy in social structures of labor, and imbalanced economic and political power. Embedded in the social system, AI technology acts as a mediator, is shaped by these root causes and in turn shapes the unethical impacts of AI in society [74, 24]. Tackling the malicious use to solve unethical issues in AI only treats the symptoms, not the disease. Similarly, the imagination that AI can fix its own issues [30] in S1 is not based on valid logic. Solving a problem requires first conducting a precise diagnosis of the root cause, and then identifying treatment that can target the diagnosis. There is no guarantee that the right treatment will always fall into the realm of the technical toolbox, and claiming AI can fix any issues is like selling a cure-all elixir [83].